

********************************************************************
UARS WEEKLY STATUS REPORT - 19 Nov 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 64% of
each day's data amounting to nearly 487 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring
scans for time periods with UARS telemetry or tape recording.
A poster presentation entitled "Retrieval of Thermospheric Molecular
Oxygen Profiles from Solar and Stellar Occultation Measurements"
(A51D-0816) authored by J Lumpe, L Floyd, M Snow, J M Picone, and G
Rottman based on SUSIM and SOLSTICE solar occultation data will be
made to the Fall AGU Meeting on Friday, December 17.
***********
PEM
***********
PEM2K:: failed on Sunday November 14, 2004. Fortunately SwRI had a
replacement available for this hardware. PEM2K:: was removed and SwRI
is currently configuring the replacement machine for PEM2K::. When
activated, the new machine will carry the PEM2K:: name and IP. The new
PEM2K:: should work transparently since the new system is also a DEC
Alpha.
HEPSA data files are version 2. Currently, HEPSA data between
1991274 and 2004199 has been reprocessed and archived at the GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004199.
PEM continues to receive NMC data files from NOAA. No problems
have been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron
data for the October and November storms of 2003 has been moved and is
now located at http://climatology.space.swri.edu/uars_fromRudy.html.
For those users of SDDAS, update your UARS Meta data and these data
times will appear with full functionality. Proton data from these
storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview
plots were produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of
2000, March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of
2001, and November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps,
array voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4
minutes, array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum
beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 06 November 2004 (Orbit 71992, GMT Day 311) through
12 November 2004 (Orbit 72097, GMT Day 317).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 312/1257 72015 POWER ON
PEM 313/1354 72031 POWER ON TO DAY ONLY
The beta angle ranged between the angles of 15.7 degrees to 32.5
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 38 degrees on 18 November 2004 (DOY 323).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.30 degrees C. Additional
battery performance data for each battery on 05 November and 12
November 2004 are provided in Appendix D.
The clock error ranged from 11.1 to -10.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 01
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 06
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Six (6) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to work on integrating the SMA
functionality. Testing was done regarding the file transfer issue
between the User Planning System and CMS to help isolate the problem.
This testing revealed that the issue may lie in the new CMS code, so
sections of the code are being review by software development.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 December 2004 (DOY 337).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 November 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
72015 312/1257 04-318 HALOE POWER ON
72018 312/1655 04-319 HALOE POWER ON - PART II
72031 313/1354 04-320 PEM POWER ON TO DAY ONLY
72037 313/2311 04-321 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
72054 315/0229 04-322 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
72061 315/1427 04-323 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72077 316/1538 04-324 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 November 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 November 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
72037 313/2327 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
72054 315/0232 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
72061 315/1429 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
72077 316/1538 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 November 2004
05 November 2004 - GMT Day 310 Beta = 12.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 -22.4/- 72.8 +5.14/+4.50 +28.0/-12.4 26.6 1.02 18.9
3 00.0/- 78.4 +3.87/+2.91 +26.8/-12.0 26.4 1.03 18.3
12 November 2004 - GMT Day 317 Beta = 32.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 05.6/- 67.2 +6.76/+5.46 +28.0/-15.2 26.1 1.04 20.3
3 -5.6/-140.0 +5.46/+4.50 +27.2/-14.8 26.1 1.04 19.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 November 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-085 FOT 71999 311/1117 37512 TDRSS 275 Dropout/Eclipse Prim
04-086 FOT 72002 312/0018 30832 Generic Late Acquisition
04-087 FOT 72009 312/0239 37516 TDRSS 275 Dropouts
04-086 FOT 72010 312/0456 30832 Generic Late Acquisition
04-086 FOT 72013 312/0934 30832 Generic Late Acquisition
04-086 FOT 72024 313/0235 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-083 FOT 71977 310/0025 N/A UARS Eclipse Server Failure
04-084 FOT 71985 310/1303 37501 TDRSS-275 dropouts
04-066 FOT See List Oct 04 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-085 FOT 71999 311/1117 37512 TDRSS 275 Dropout/Eclipse Prim
04-086 FOT See List Nov 04 30832 Generic Late Acquisition
04-087 FOT 72009 312/0239 37516 TDRSS 275 Dropouts